an angle is 3 times more than its complement find each angle
Found 2 solutions by erica65404, Alan3354:
Answer by erica65404(394)

complementary angles add up to 90
x+y=90
3x=y
plug into equation
x+3x=90
4x=90
x=22.5
now plug this into original equation
22.5+y=90
y=67.5
The angles are 22.5 and 67.5

Answer by Alan3354(69443)

an angle is 3 times more than its complement find each angle
-----------
1 time more = 2x as much
--> 3 times more = 4 times as much
A = 4*(90 - A)
A = 360 - 4A
5A = 360
A = 72 degs